two loops of wire of radii 1cm and 2cm resp.are placed coaxialy .current i passes through both in same direction initialyand then the the direction of current is reversed . what is the ratio of magnetic induction at centre in first and second case

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear darshan joshi

magnetic field at the center of loop of radius a is

B = μ i/2a

first case

B=B1 +B2

  = μi/2*1    + μ i/2*2

  = μi/2 [ 3/2]


second case

B' = B1 -B2

   =B1 +B2

  = μi/2*1    - μ i/2*2

  = μi/2 [ 1/2]

so B/B'  = 3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin